OpenStudy (anonymous):

Can someone check if i am doing this right?? James has decided to start saving for a cruise over the summer. His money is currently in the local bank, modeled by the function s(x) = 102. He is able to do work around the neighborhood to earn extra money and that function would be a(x) = 8(x + 2), where x is measured in hours. Explain to James how he can create a function that combines the two, and describe any simplification that can be done.

OpenStudy (anonymous):

My answer is : To combine both of the functions you could add then S(x) + a(x) = 102+ 8(x+2) when you solve it you woud get 102+ 8(x+2) 102 + 8x + 16 118 + 8x= 0 -118 -118 8x= -118 __ ___ 8 8 x= 14.75

OpenStudy (anonymous):

u cant solve bc there isnt enough info

OpenStudy (anonymous):

u dont know what it adds to but u could do this

OpenStudy (anonymous):

@Lala3712

OpenStudy (anonymous):

Ohh Okay. So then would I just leave out the part where i solved it and keep the other part the just words part?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

fan and medal plus mention me for help @Lala3712

OpenStudy (anonymous):

Thank you